which number is the biggest? 5/7 ot 6/9 . calculate the answer using two methods. I got 5/7 as the largest

Answer by Alan3354(69443)

which number is the biggest? 5/7 ot 6/9 . calculate the answer using two methods. I got 5/7 as the largest
----------------
One way is to convert to decimal
5/7 = 0.7142857...
6/9 = 0.6666666...
-----------
A 2nd way is to find the LCD
5/7 = 15/21
6/9 = 14/21
Either way, 5/7 is bigger.